[figure: 9.7 cm]

SUPERCONDUCTIVITY IN DISORDERED SYSTEMS

G. Litak^{a}, T. Paiva^{b}, R.T. Scalettar^{c}, C. Huscroft^{d}, R.R dos Santos^{b} aTechnical University of Lublin, Nadbystrzycka 36, PL–20-618 Lublin, Poland bInstituto de Fisica, Universidade Federal dos Rio de Janeiro, Cx.P. 68 528, 21945-970 Rio de Janeiro RJ, Brazil cPhysics Department, University of California, Davis, CA 95616 dHewlett-Packard Company, 8000 Foothills Boulevard, Roseville, CA 95747

We study the effect of non-magnetic impurities on superconductivity in the negative U Hubbard model on a two dimensional square lattice. Two kinds of disorder are discussed and their influences on superconductivity are compared. The first kind of disorder deals with random site energy ϵ_{i}. The second one is connected with random interaction U_{i}, which can be −|U| and 0 with probability c and 1 − c depending on the lattice site i. Using the quantum Monte Carlo approach we find the critical concentration of impurity sites which destroys superconductivity.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : oral

Corresponding author : G. Litak

Address for correspondence : Technical University of Lublin, Nadbystrzycka 36, PL–20-618 Lublin, Poland

Email address : litak@archimedes.pol.lublin.pl